Two payments of $49 000 each must be made 3 year and 5 year from now. If money can earn 4.9% compounded monthly, what single payment 5 years from now would be equivalent to the two scheduled payments?
Answer #1

Present value now =Future value/(1+Rate)^number of periods

= 49000/(1+4.9%/12)^(3*12) + 49000/(1+4.9%/12)^(5*12)

= 42314.07+ 38371.66

= 80685.73

Value of the payment 5 years from now = PV*(1+rate)^n

= 80685.73*(1+4.9%/12)^(5*12)=

103034.40